Southport Company is considering the purchase of a piece of equipment that costs $100,000. The equipment would be depreciated on a straight-line basis to its expected salvage value of $10,000 over its 10-year useful life. Assuming a tax rate of 40%, what is the annual amount of the depreciation tax shield provided by this investment?

Multiple Choice

  • $4,000

  • $9,000

  • $3,600

  • None of these answers is correct.
Answer #1

Annual depreciation=(Cost-salvage value)/Useful Life

=(100,000-10000)/10=$9000/year

Hence annual depreciation tax shield=Annual depreciation*Tax rate

=(9000*40%)

=$3600